JRN Enterprises just announced that it plans to cut its dividend from $2.75 to $1.40 per share and use the extra funds to expands its operations. Prior to this announcement, JRN's dividends were expected to grow at 6% per year and urn's stock was trading at $24.00 per share, With the new expansions, JRN's dividends are expected to grow at 12% per year indefinitely. Assuming that JRN's risk is unchanged by the expansion, the value of a share of JRN after the announcement is closest to:

a. $50.38

b. $25.65

c. $12.22

d. $24.00
Answer #1

First, we need to find the required return, rE

rE = [D1 / P0] + g

= [$2.75 / $24] + 0.06 = 0.1146 + 0.06 = 0.1746, or 17.46%

Now, we can find the new stock price;

P0 = D1 / [rE - g]

= $1.40 / [0.1746 - 0.12] = $1.40 / 0.0546 = $25.65

Hence, Option "b" is correct.